### v2

# Dynamics of a compressible gas injected into a confined porous layer

Abstract
Underground gas storage is a critical technology in global efforts to mitigate climate change. In particular, hydrogen storage offers a promising solution for integrating renewable energy into the power grid. When injected into the subsurface, hydrogen’s low viscosity compared to the resident brine causes a bubble of hydrogen trapped beneath caprock to spread rapidly into an aquifer through release of a thin gas layer above the brine, complicating recovery. In long aquifers, the large viscous pressure drop between source and outlet induces significant pressure variations, potentially leading to substantial density changes in the injected gas. To examine the role of gas compressibility in the spreading dynamics, we use long-wave theory to derive coupled nonlinear evolution equations for the gas pressure and gas/liquid interface height, focusing on the limit of long domains, weak gas compressibility and low gas/liquid viscosity ratio. Simulations are supplemented with a comprehensive asymptotic analysis of parameter regimes. Unlike the near-incompressible limit, in which gas spreading rates are dictated by the source strength and viscosity ratio, and compressive effects are transient, we show how compression of the main gas bubble can generate dynamic pressure changes that are coupled to those in the thin gas layer that spreads over the liquid, with compressive effects having a sustained influence along the layer. This coupling allows compressibility to reduce spreading rates and gas pressures. We characterise this behaviour via a set of low-order models that reveal dominant scalings, highlighting the role of compressibility in mediating the evolution of the gas layer.
1 Introduction
Underground gas storage plays a crucial role in addressing anthropogenic climate change, both by sequestering greenhouse gas emissions prior to atmospheric release and by enabling the transition to cleaner energy sources. The first approach is well-established through carbon capture and storage, where large quantities of CO2 are injected underground for long-term containment. More recently, underground hydrogen storage has gained significant attention for its potential to enable large-scale integration of renewable energy into the power grid (heinemann_enabling_2021; zivar_underground_2021; muhammed_review_2022; lord_geologic_2014; tarkowski_underground_2019). The concept is to use surplus renewable energy generated during the summer months to produce hydrogen via electrolysis, store it in underground formations, and subsequently recover it during the winter when demand exceeds supply. Hydrogen is a particularly attractive candidate for this large-scale energy storage owing to its high energy density and the fact that its combustion produces minimal harmful emissions (zivar_underground_2021). Proposed underground storage sites for gas include depleted reservoirs, aquifers, and salt caverns, all of which are typically saturated with brine. Unlike CO2 storage, where the gas is intended for permanent sequestration, hydrogen storage requires recoverability, making accurate predictions of gas plume evolution crucial. Due to its exceptionally low viscosity relative to brine, injected hydrogen may spread rapidly along the top of the brine within an aquifer, forming a thin gas layer that extends far from the injection site (hagemann_mathematical_2015). This behaviour not only reduces storage efficiency but also increases the risk of hydrogen loss.
The injected layer of hydrogen, through its displacement of heavier brine, is a form of viscous gravity current (zheng2022). Much of the research on viscous gravity currents has focused on incompressible flows, where the interaction between buoyancy and viscosity governs the rate and extent of spreading. In the context of hydrogen storage, the low viscosity of the injected gas, coupled with buoyancy effects, plays a critical role in determining how far and how quickly the hydrogen spreads underground. Early studies primarily addressed gravity-driven spreading, where a dense fluid advances along a rigid boundary within an infinitely deep porous medium (barenblatt_unsteady_1952; huppert_gravity-driven_1995; lyle_axisymmetric_2005). However, when a fluid is actively injected into a confined porous layer, the resulting pressure gradient, caused by the pressure drop between the injection site and the far field, significantly alters the flow dynamics.
huppert_gravity-driven_1995 examined the exchange of two confined incompressible fluids with differing densities and pressures, but equal viscosities, between two aquifers. By deriving a similarity solution, they showed that buoyancy transports the denser fluid along the lower boundary and the lighter fluid along the upper boundary, while an imposed pressure gradient drives a net flow. Subsequent research expanded on this framework to incorporate fluids with different viscosities, highlighting how the viscosity contrast influences spreading behaviour (nordbotten_similarity_2006; pegler_fluid_2014; zheng_flow_2015). nordbotten_similarity_2006 generalized Huppert & Woods’ (huppert_gravity-driven_1995) study to investigate the effect of viscosity contrasts in an axisymmetric geometry. pegler_fluid_2014 conducted a comprehensive study of flow behaviour in a two-dimensional channel. Their findings revealed that, at early times, spreading is primarily buoyancy-driven and can be approximated by the classical porous-medium equation of barenblatt_unsteady_1952. At later times, low-viscosity injected fluid forms a thin layer along the upper boundary, with buoyancy becoming less dominant compared to viscous effects. By deriving a large-time similarity solution, they demonstrated that the movement of the upper contact line is controlled by the strength of the source. In this case, since the flow is incompressible, the length of the channel determines the pressure needed to drive the motion but does not affect the spreading dynamics. These theoretical predictions were validated through laboratory experiments in which freshwater was injected into a saltwater-saturated porous medium, confirming that the similarity solution accurately describes the interface evolution.
When the injected gas is compressible, however, the dynamics change: the source can in principle increase the gas pressure without advancing the gas-liquid interface, and the subsequent propagation of the upper contact line may be influenced by the length of the channel. In long channels, the viscous pressure drop between source and outlet may lead to spatial variations in pressure, which could in turn lead to noticeable variations in gas density along the channel. These variations could affect the mass flow rate and, consequently, influence the rate of gas spreading. Despite the potential for compressibility to significantly alter the flow dynamics, its effects in this context remain under-explored. While some studies have addressed compressibility, most have focused on fluid-structure interactions, particularly the impact of pressure buildup on the rock matrix integrity (mathias_approximate_2009) and fluid migration in neighboring reservoirs (jenkins_impact_2019). A common approach to coupling the mechanics of the pore structure with pressure build-up during injection is to introduce rock compressibility, defined as $c_{r}=({1}/{\phi}){\mathrm{d}\phi}/{\mathrm{d}p^{*}}$, where $\phi$ is the porosity and $p^{*}$ is the pressure. This parameter quantifies how the pore space deforms in response to pressure changes and is valid under conditions of constant vertical stress and negligible lateral strain (jenkins_impact_2019).
mathias_approximate_2009 examined pressure build-up during gas injection into an axisymmetric channel of infinite extent. Using matched asymptotics, they derived similarity solutions for pressure evolution in the case where the compressibilities of the gas and resident liquid are comparable (relative to rock and water), and where the interface propagates much more slowly than the diffusive pressure front. This work was later extended by mathias_pressure_2011 to finite-length channels, offering further insight into pressure evolution under more realistic conditions. jenkins_impact_2019 investigated pressure dissipation in a layered aquifer system, where multiple confined channels are stacked vertically. They analyzed how water leakage from the injection aquifer to the surrounding aquifers influences the pressure field. In their model, CO2 displaces resident water in the injection channel, and weak vertical flow allows liquid to escape through the confining boundaries, relieving pressure within the system. By incorporating the compressibility of both gas and liquid through a linearized equation of state, they found that pressure reduction due to water escape slows the gas plume’s spreading rate. In contrast to these studies, cuttle_compression-driven_2023 examined how compressibility interacts with surface tension and viscous forces in the liquid, to influence viscous fingering instabilities, when a compressible gas is injected into a Hele-Shaw cell filled with water. By neglecting pressure variations in the gas due to viscous effects, they used Boyle’s law to couple the uniform pressure in the gas, which drives displacement, to the viscous effects in the liquid and the capillary forces at the interface. Their results showed that gas compressibility delays the onset of viscous fingering and reduces its severity, with the severity quantified by the isoperimetric ratio (which compares the length of the interface to the area it occupies). Additionally, compressibility was found to increase the breakthrough time, i.e. the moment when the interface reaches the outlet.
In this paper, we extend the framework of pegler_fluid_2014 for injection of a gas into a confined porous channel containing brine (figure 1) to investigate how gas compressibility interacts with buoyancy and viscous effects in both fluids to determine the spreading dynamics. Using a long-wave approximation, we derive two coupled nonlinear evolution equations governing the gas/liquid interface height and the gas pressure field. These are mass conservation equations, with the gas pressure serving as a proxy for gas density via the equation of state. The viscous pressure drop ahead of the upper contact line is captured by a boundary condition, while two kinematic conditions determine the motion of the contact lines. The model admits several distinguished asymptotic limits that reveal the roles of different mechanisms across parameter space; we derive a set of reduced models that capture dominant balances. In particular, we estimate key quantities of practical interest, including the pressure scale, pressure rise time, and the breakthrough time. In §2, we present the governing equations and describe the scaling argument used to derive the long-wave model. The system is characterized by three primary dimensionless parameters, and we simplify the model in various sub-limits, identifying relevant scaling relationships. In §LABEL:sec:results, we present numerical solutions of the full model, illustrating how the key features of compressible flow emerge and comparing them with predictions from asymptotic analysis. Finally, in §LABEL:sec:discussion, we revisit the dominant balances in each regime, interpret the underlying physical mechanisms, and discuss the broader implications for hydrogen storage and related gas-injection technologies.
Figure 1: Schematic showing the displacement of an ambient liquid in a porous medium (regions IIw and III) due to the injection of a compressible gas (occupying regions I and IIg) from a line source at $\boldsymbol{x}^{*}=\boldsymbol{x}_{0}^{*}$. The interface separating the fluids is sharp and located at ${y}^{*}={F}^{*}({x}^{*},{t}^{*})$. The pressure in the ambient brine is hydrostatic at the outlet at $x^{*}=L$.
2 Model formulation
2.1 Governing equations
We consider a two-dimensional model of fluid displacement driven by the injection of a compressible gas into a planar horizontally-confined porous medium with depth $H$ and length $L$, as illustrated in figure 1. Initially, the medium is predominantly saturated with a liquid, with only a small amount of gas present. The length of the region initially occupied by gas is denoted by $L_{g0}$; the aspect ratio of this region is assumed small, i.e. $H/L_{g0}\ll 1$. (We do not seek here to model the initial formation of the gas bubble.) The permeability of the medium is taken to be uniform, with value $k_{0}$.
At time $t^{*}=0$, gas is injected into the channel from a line source located at $\boldsymbol{x}^{*}~\equiv(x^{*},y^{*})=\boldsymbol{x}^{*}_{0}$. Variations in the injection rate are governed by the function $q\,\mathcal{Q}(\omega{t^{*}})$, where $1/\omega$ represents the timescale of injection, $q$ denotes the source strength per unit width of the channel, and $\mathcal{Q}$ is a dimensionless function. The pressure at the outlet ($x^{*}=L$) is taken to be hydrostatic with a baseline value of $p_{g0}$. Using a sharp-interface model, disregarding capillary or miscibility effects, the fluids are segregated into two distinct phases and the domain is divided into four regions (figure 1): the gas occupies regions I and IIg, while the liquid occupies regions IIw and III. We use the subscripts $g$ and $w$ to represent quantities in the gas and liquid (water) phase respectively. We revisit all modelling assumptions in §4.
We denote the height of the gas-liquid interface by $y^{*}=F^{*}(x^{*},t^{*})$ and the points of contact with the upper and lower boundaries by $X^{*}_{u}({t^{*}})$ and $X^{*}_{l}({t^{*}})$ respectively. For later convenience, we extend the definition of $F^{*}$ to the whole domain such that
$F^{*}(x^{*},t^{*})=\begin{cases}0,&0\leq x^{*}\leq X^{*}_{l}(t^{*}),\\ H,&X^{*}_{u}(t^{*})\leq x^{*}\leq L.\end{cases}$ (1)
[…]
The liquid is assumed to be incompressible with constant density ${\rho}_{w}$. The initial gas and liquid volumes are assumed to be in thermal equilibrium with the surrounding pore matrix. Gas compression is treated as isothermal, an assumption that is valid when the thermal inertia of the solid pore matrix is much greater than that of the gas (kushnir_temperature_2012) At a fixed temperature, we define $\rho_{g0}$ to be the equilibrium gas density when at the outlet pressure $p_{g0}$. Then, using $p_{g0}$ and $\rho_{g0}$ as a reference pressure and a reference density, we write the equation of state for the gas (relating density $\rho^{*}_{g}$ to pressure $p^{*}_{g}$) and the equilibrium speed of sound in the gas ($c$) as
$p^{*}_{g}=p_{g0}\mathcal{P}\left(\frac{\rho^{*}_{g}}{\rho_{g0}}\right),\qquad c^{2}\equiv\left.\frac{\partial{p^{*}_{g}}}{\partial\rho^{*}_{g}}\right|_{\rho_{g0}}=\frac{p_{g0}}{\rho_{g0}}\mathcal{P}^{\prime}(1),$ (2)
respectively, for a suitable dimensionless function $\mathcal{P}$, for which $\mathcal{P}(1)=1$ and $\mathcal{P}^{\prime}(1)>0$. For hydrogen, $\mathcal{P}^{\prime}(1)\geq 1$ and $\mathcal{P}^{\prime\prime}(1)\geq 0$ at temperatures and pressures of interest (figure 2bb), so that $c^{2}$ rises with increasing pressure, although this may not be the case for gases such as CO2 (figure 2aa).
The fluid velocities $\boldsymbol{u}_{i}^{*}=(u^{*}_{i},v^{*}_{i})$ in both phases are then given by
$\boldsymbol{u}^{*}_{i}=-\frac{k_{0}}{\mu_{i}}\>(\boldsymbol{\nabla}^{*}p^{*}_{i}-\rho^{*}_{i}\boldsymbol{g}),\qquad(i=w,g),$ (3)
where $\boldsymbol{g}$ is the acceleration due to gravity, $\rho_{w}^{*}\equiv\rho_{w}$ and $\mu_{g}$ and $\mu_{w}$ are gas and liquid viscosities, assumed constant. The pore-scale Reynolds number, $Re=qd/(H\mu_{g})$, is assumed sufficiently small to neglect inertia, where $d$ is a typical pore diameter. Taking $H=10\,$\mathrm{m}$\$, $\mu_{g}\approx 10^{-6}\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-1}\,\mathrm{s}^{-1}$\$ and $d\approx 10\,\mu$\mathrm{m}$\$ (hashemi_pore-scale_2021), the condition $Re\leq 1$ suggests that inertia can be neglected for injection rates up to $q\approx 1\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-1}\,\mathrm{s}^{-1}$\$.
The continuity equations for the gas and liquid phases are
$\rho^{*}_{g,t^{*}}+\boldsymbol{\nabla}^{*}\cdot(\rho^{*}_{g}\boldsymbol{u}^{*}_{g})=q\,\mathcal{Q}(\omega t^{*})\,\delta^{*}(\boldsymbol{x}^{*}-\boldsymbol{x}_{0}^{*}),\quad\boldsymbol{\nabla}^{*}\cdot\boldsymbol{u}_{w}^{*}=0,$ (2.4a,b)
respectively, where $\delta^{*}(\boldsymbol{x}^{*})$ is the Dirac delta function. The source at $\boldsymbol{x}^{*}=\boldsymbol{x}_{0}^{*}$ lies within a distance of order $H$ of an impermeable boundary at $x^{*}=0$. Equations (2), (3) and (2.4aa) form the governing equations for the gas and are applied in regions I and IIg; (3) and (2.4bb) apply in regions IIw and III. The kinematic condition
$F^{*}_{t^{*}}+u^{*}_{i}F^{*}_{x^{*}}=v^{*}_{i},\quad\text{at}\quad y^{*}=F^{*},\quad(i=w,g;~~0\leq x^{*}\leq L),$ (5)
relates the velocity of the interface to the fluid velocities for $X^{*}_{l}<x^{*}<X^{*}_{u}$. Subscripts $x^{*}$ and $t^{*}$ denote partial derivatives. No-penetration at the walls implies $u^{*}_{g}=0$ at $x^{*}=0$, $v^{*}_{i}=0$ at $y^{*}=H$ and $v^{*}_{i}=0$ at $y^{*}=0$. The relative strength of buoyancy to capillary forces can be characterized by a Bond number. We adopt the definition of the Bond number from GOLDING_2011 and zheng_self-similar_2019, which measures the relative magnitude of buoyancy to the capillary entry pressure $p_{e}$ in the largest pores. Using $p_{e}=\gamma/k_{0}^{1/2}$ (zheng_self-similar_2019), the Bond number can be expressed as $Bo=\Delta\rho gHk_{0}^{1/2}/\gamma$, where $\Delta\rho=\rho_{w}-\rho_{g0}\approx 10^{3}\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-3}$\$ is the density difference between the fluids, $g$ is gravitational acceleration, $H\approx 10$$50\,$\mathrm{m}$\$ is a characteristic aquifer height, $k_{0}\approx 10^{-14}$$10^{-12}\,$\mathrm{m}^{2}$\$ is the permeability, and $\gamma\approx 7\times 10^{-2}\,$\mathrm{N}\,\mathrm{m}^{-1}$\$ is the hydrogen–liquid interfacial tension (zivar_underground_2021). An order-of-magnitude estimate gives $Bo\sim 0.1\,\text{--}\,10$, indicating that capillary effects may be important in shallow, low-permeability reservoirs. A detailed treatment of the capillary fringe is beyond the scope of this study, although we discuss its potential influence in §4. Accordingly, we assume a sharp interface and take the pressure to be continuous across it: